OpenStudy (anonymous):

A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@eliassaab @ganeshie8 @RadEn

OpenStudy (anonymous):

8x + 5y = 9 3x + 2y = 4 First solve the two equations, you find x =-2 and y=5 Multiply the first equation by 2 and add it to second, you obtain 8x + 5 y =9 19 x + 12 y = 22 verify that x =-2 and y =5 are still solutions of the new system.